# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ Acknowledgements
Table A.1: Key terminologies for LLM-based agentic AI systems.
@@
Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
@@
Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
@@
Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
@@
Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
@@
ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
@@
Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
@@
Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
@@
Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
@@
Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
Figure A.1: Traditional AI Agent vs. Agentic AI.

### v5

@@ Acknowledgments
Author Contributions
S.R. and R.S. contributed equally to this work. S.R.: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing, Revision R.S: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing M.K.: Supervision, Detailed review, Structuring, Writing. C.E.: Detailed review, Writing, Re-writing, Supervision, Revision. All authors have read and approved the final manuscript.
@@
Table A.1: List of abbreviations used in this paper.
Abbreviation Description
AI Artificial Intelligence
AMAS Agentic Multi-Agent Systems
CoT Chain-of-Thought
CSS Component Synergy Score
DP Differential Privacy
FHE Fully Homomorphic Encryption
HITL Human-in-the-Loop
HE Homomorphic Encryption
IAM Identity and Access Management
LLM Large Language Model
MLOps Machine Learning Operations
ModelOps Model Lifecycle Operations, Monitoring, and Governance
MPC Secure Multi-Party Computation
PII Personally Identifiable Information
RAG Retrieval-Augmented Generation
RBAC Role-Based Access Control
RMF Risk Management Framework
TEEs Trusted Execution Environments
TRiSM Trust, Risk, and Security Management
TUE Tool Utilization Efficacy
XAI Explainable Artificial Intelligence
Figure A.1: Traditional AI Agent vs. Agentic AI.
Table A.2: Key terminologies for LLM‐based agentic AI systems.
@@
Agentic AI System A multi-agent artificial intelligence system powered by large language models (LLMs), in which multiple autonomous agents coordinate, communicate, and reason collectively to plan and execute tasks over extended horizons. Such systems typically incorporate persistent memory, tool use, and dynamic role assignment to support adaptive, goal-driven behavior.
Autonomy Model The internal decision-making mechanism that governs how an agent selects goals, plans actions, and executes tasks without continuous human intervention, often leveraging planning, reasoning traces, feedback signals, and environmental context.
@@
Chain of Thought (CoT) A prompting and reasoning paradigm in which an LLM explicitly generates intermediate reasoning steps before producing a final output, enabling multi-step problem solving, improved transparency, and more controllable agent behavior.
Counterfactual Analysis An interpretability and diagnostic technique that evaluates how changes to inputs, agent roles, or internal states would alter system outcomes, helping reveal causal dependencies, failure modes, and contribution disparities across agents.
@@
Explainability The ability of an AI system or individual agent to provide human-understandable rationales for its decisions, actions, or outputs, often supported by techniques such as feature attribution, decision provenance, or structured reasoning traces.
@@
Foundation Model (LLM) A large pretrained language model (e.g., GPT-4, LLaMA) that serves as the core reasoning and generation component of an agent, providing linguistic competence, contextual understanding, tool invocation, and general-purpose reasoning capabilities.
Shared Memory (Persistent Memory) A centralized or distributed memory substrate, commonly implemented using vector databases or structured logs, that enables agents to store, retrieve, and share contextual information across time, supporting long-term coherence and coordination.
@@
ModelOps A set of operational practices for managing models and agent prompts throughout their lifecycle, including versioning, deployment, monitoring, auditing, and retirement, with an emphasis on reliability, reproducibility, and governance.
Application Security Protective mechanisms and design practices that safeguard agentic systems against adversarial behaviors such as prompt injection, identity spoofing, privilege escalation, and cross-agent manipulation.
@@
Model Privacy Privacy-preserving techniques that ensure sensitive data remains protected during model training, inference, and inter-agent communication, including differential privacy, secure computation, and trusted execution environments.
Prompt Injection A class of security attacks in which maliciously crafted inputs introduce hidden or indirect instructions that override or corrupt an agent’s intended behavior, potentially propagating harmful actions across agent interactions.
Retrieval-Augmented Generation (RAG) An architectural approach in which agents dynamically retrieve relevant external knowledge (e.g., documents or facts) from a datastore and condition LLM generation on this retrieved context to improve factuality and grounding.
Role-Specialized Coordination A coordination strategy where agents are assigned distinct functional roles (e.g., planner, verifier, executor) and interact via structured communication protocols to collectively solve complex tasks more efficiently and robustly.
@@
Decision Provenance Graph A structured, graph-based representation that records data flows, reasoning steps, and inter-agent interactions, enabling post-hoc auditing, accountability analysis, and system-level interpretability.
Tool-Use Interface The mechanism that allows agents to issue structured commands to external tools, APIs, or execution environments and integrate returned results into subsequent reasoning and decision-making steps.
@@
Trust Score A composite quantitative indicator of an agent’s reliability and alignment, typically combining task performance, safety compliance, behavioral consistency, and calibration of expressed confidence over time.
@@
Composite Metric An aggregated evaluation measure that combines multiple dimensions—such as trustworthiness, explainability, coordination quality, and user-centered performance—to benchmark and compare agentic AI systems holistically.